If spherical metal ball has a mass of 1.00E3 g and a density of 10.9 g/mL. What is its diameter (in cm)?
The volume of a sphere can be calculated using V= (4/3)πr3

Answer by Alan3354(21581)

If spherical metal ball has a mass of 1.00E3 g and a density of 10.9 g/mL. What is its diameter (in cm)?
---------------
Mass = Volume x Density
1000 = V*10.9
V = 1000/10.9 = (4/3)*PI*r^3
r^3 = 1000/(10.9*(4/3)*PI
r^3 = 3000/(10.9*4*PI)
r^3 = 21.902
r = 2.79787... cm
D = r*2 = 5.55974... cm